# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v2

@@ Abstract
The modulo analog-to-digital converter (ADC) is a promising solution to resolve the limited dynamic range (DR) issue of conventional ADCs and achieve an enhanced digital resolution given a fixed quantization bit budget. However, a modulo ADC requires an unfolding scheme to correct the nonlinear distortion introduced by the modulo operation. This paper presents a sliding discrete Fourier Transform (DFT)-based method for fast signal reconstruction given the modulo ADC output sequence and a 1-bit folding information sequence. In contrast to existing DFT-based signal recovery techniques for modulo ADCs, our proposed sliding DFT method reduces the required observation time and minimizes the spectral leakage effects via proper choice of window function parameters. A mean squared error (MSE) performance guarantee is established for the proposed signal recovery algorithm. More precisely, we derive sufficient conditions for the oversampling factor ($\mathrm{OF}$) and the number of quantization bits ($b$) to obtain a specific MSE performance. Our numerical results demonstrate that modulo ADCs equipped with our proposed recovery method can outperform conventional ADCs without modulo for $\mathrm{OF}\geq 4$ and $b\geq 4$. The impact of spectral leakage on the MSE performance of the proposed sliding DFT recovery method is also quantified.
@@ I Introduction
A promising solution to address the DR bottleneck of ADCs is the modulo sampling framework [6, 7]. The idea in modulo sampling is to apply a nonlinear folding operation called modulo to the input signal before the ADC. This modulo preprocessing stage ensures that the signal range is confined within the DR of the ADC. Moreover, for a given amplitude quantization bit budget, modulo ADCs can achieve better digital resolution compared to conventional ADCs as demonstrated in prior works [8, 9].
@@
Since modulo operation intentionally introduces nonlinear distortion to the signal, several studies on modulo sampling have focused on the development of robust recovery techniques to correct the distortion induced by the modulo operation. Reconstruction techniques for modulo sampling include higher-order differences-based approaches [7], discrete Fourier transform (DFT)-based techniques [10, 11, 12, 13, 14, 15], prediction-based filtering [16, 17], iterative methods [18], sparse signal modulo recovery [19, 20, 21, 22], multi-channel modulo structures [23, 24, 25, 26], and thresholding-based schemes[27, 28, 29, 30, 31]. There are also research efforts geared toward the practical hardware implementation of modulo ADCs [10, 32, 33, 34, 35, 36] and their potential applications in radars, communication systems, imaging, and electroencephalogram (EEG) recovery [37, 38, 39, 40, 12, 41, 42, 43, 25, 44, 45].
@@
Another important research direction in modulo ADC is the establishment of theoretical performance guarantees to reconstruct the unfolded signal from modulo samples. In [7], it is shown that the higher-order differences-based reconstruction approach can perfectly reconstruct a finite-energy bandlimited (BL) signal from its modulo samples if the sampling rate used is at least $2\pi e$ ($\approx 17.07$) times the Nyquist rate. This sufficient condition for the sampling rate increases significantly in the presence of bounded noise [7, Theorem 3]. For finite-energy BL signals, perfect reconstruction is possible for any sampling rate approaching the Nyquist rate from above by using a prediction filter with sufficiently long filter length [16]. This approach does not need any side-information such as knowledge of unfolded samples and auto-correlation function. However, the proposed prediction filter-based approach did not take into account the impact of finite quantization. Theoretical guarantees for Fourier-based reconstruction had been established in [10]. However, this approach suffers from spectral leakage and may require long observation windows. To avoid spectral leakage, recovery methods based on time-domain thresholding can also be used. Such methods work for generalized modulo sampling models which consider the modulo hysterisis and the folding transients. However, recovery guarantees of thresholding-based recovery fail under ideal modulo nonlinearities. This is explained in Section III-E. A recent work on modulo ADCs with 1-bit side information [14] established a mean squared error (MSE) performance guarantee when the oversampling factor (OF) is above three and the number of quantization bits is above three. The 1-bit side information identifies the indices in which folding occured. However, the proofs neglected the spectral leakage effects and the contribution of quantization noise at certain time indices.
@@
In this work, we propose a sliding DFT approach to construct the original input samples from the modulo ADC output samples and an additional 1-bit folding information signal. The recovery procedure uses short-time Fourier transform (STFT) to facilitate successive frame-based unfolding of the modulo samples. To significantly reduce the impact of spectral leakage in STFT computation, a window function with smooth transitions at the edges is multiplied to the frame. A MSE performance guarantee for the proposed recovery method is also established and is compared with the simulated MSE results.
@@
  • We compare the MSE of the modulo ADC equipped with our proposed recovery procedure to that of the conventional ADC, i.e., no modulo operation. Under the established sufficient conditions, we show that the MSE of the modulo ADC which uses the sliding DFT recovery algorithm is in $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ when spectral leakage is neglected and is in $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ otherwise. The parameter $\mathrm{OF}$ is the oversampling factor to be defined in Section II. Meanwhile, the MSE of a conventional ADC is in $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$. Numerical results are presented to demonstrate these behaviors.
Notation: The following notations are used throughout the paper. The sets of real numbers, integers, and natural numbers are denoted as $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{N}$, respectively. When referring to discrete-time signals, the notation $z[n]$ is used to signify $z(nT_{\mathrm{s}})$, assuming that the sampling period $T_{\mathrm{s}}$ is evident from the context. The first-order difference of a discrete-time signal $x[n]$ is denoted as $\underline{x}[n]=x[n]-x[n-1]$, where $x[-1]=0$ unless stated otherwise. Vectors and matrices are written in bold format (e.g., $\mathbf{z}$, $\mathbf{A}$) while sets are written in calligraphic format (e.g., $\mathcal{S}$). The cardinality of a set $\mathcal{S}$ is written as $|\mathcal{S}|$. If $\mathbf{x}$ is a vectorized form of some discrete-time signal $x[n]$, then $\underline{\mathbf{x}}$ is the vectorized form of $\underline{x}[n]$. The $\ell_{p}$-norm of a vector $\mathbf{x}$ is written as $\|\mathbf{x}\|_{p}$. The set of values of the discrete-time signal $x[n]$ for $n\in\mathcal{S}$ is denoted as $x_{\mathcal{S}}[n]$. The Moore-Penrose inverse of a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ is denoted by $\mathbf{A}^{\dagger}$. We use $\omega$ and $\Omega$ to denote the frequencies for the contininuous-time Fourier Transform (CTFT) and discrete-time Fourier Transform (DTFT) spectrum, respectively. We also use the standard Big-O notation $\mathcal{O}(\cdot)$ to describe asymptotic growth rates in this paper.
@@ II-A Signal Acquisition Process
We consider the modulo ADC system shown in Figure 1. The input $f(t)$ is a bandlimited signal with (angular) frequency support $\left[-\frac{\omega_{m}}{2},\;+\frac{\omega_{m}}{2}\right]$ and zero mean. This signal is folded by the modulo operator block to produce the modulo signal $f_{\lambda^{\prime}}(t)$. The operation behind the modulo folding mechanism is shown in Figure 2. This feedback structure for the folding operation is adopted from [34]. Comparator 1 triggers a positive value whenever $f_{\lambda^{\prime}}(t)$ crosses $+\lambda$ from below. Similarly, comparator 2 triggers a negative value whenever $f_{\lambda^{\prime}}(t)$ crosses $-\lambda$ from above. These trigger signals are fed to a discrete voltage generator (DVG). The voltage level of the DVG is increased (resp. decreased) by $2\lambda^{\prime}$ whenever it receives a positive (resp. negative) trigger signal. A detailed circuit-level implementation of the DVG is presented in [34]. Mathematically, the folded signal can be expressed as
@@
where $\lambda^{\prime}\in(0,\|f(t)\|_{\infty})$ is the modulo threshold. The $\ell_{\infty}$-norm of a waveform is its maximum amplitude. Throughout this paper, we assume ideal modulo nonlinearities. That is, there are no modulo hysterisis and folding transients like in the generalized modulo sampling model presented in [28].
To digitize the modulo signal, $f_{\lambda^{\prime}}(t)$ is first sampled at every $T_{s}$ seconds. The angular sampling rate is $\omega_{s}=\frac{2\pi}{T_{s}}=\mathrm{OF}\times\omega_{m}$, where $\mathrm{OF}\geq 1$ is the oversampling factor. We also define $\rho=\frac{1}{\mathrm{OF}}$. Prior to quantization, a (non-subtractive) dither sequence $d[n]$ is added to $f_{\lambda^{\prime}}[n]$. Samples of the dither sequence are drawn i.i.d. from a triangle distribution with amplitude support $\big{(}-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\big{]}$. The rationale for using triangle dither in the modulo ADC problem setup is explained in Section III-B. The resulting signal is then forwarded to the $b$-bit uniform scalar quantizer $\mathcal{Q}_{b}(\cdot)$ with range $[-\lambda,+\lambda]$. Consequently, each quantization bin has width $\frac{2\lambda}{2^{b}}$. To prevent the quantizer from being overloaded, the quantizer DR is set to $\lambda=\frac{2^{b}\lambda^{\prime}}{2^{b}-2}$. The quantizer output can be written as
$f_{\lambda^{\prime},\mathrm{q}}[n]=$ $\mathcal{Q}_{b}(f_{\lambda^{\prime}}[n]+d[n])$
$=$ $f[n]+z[n]+\epsilon[n],$ (2)
@@
where $\tau_{k}$ is the location of the level crossings, $g_{1}(t)$ is the pulse shape of the trigger signals, $g_{2}(t)$ is the sampling kernel, and $g(t)=(g_{2}\ast g_{1})(t)$. To simplify the analysis, $g(t)$ is assumed to be a rectangular pulse with width $T_{s}$ and unity amplitude. The signal $c(t)$ is sampled at rate $\omega_{s}$ and then fed to a 1-bit comparator to produce $c[n]$. In essence, $c[n]$ provides information if the signal crossed either $+\lambda^{\prime}$ or $-\lambda^{\prime}$ in the interval $(nT_{\mathrm{s}},(n+1)T_{\mathrm{s}}]$. We note that even though our modulo ADC generates two channel streams, our setup is different from the multichannel modulo sampling structures presented in [23, 24, 25]. The multiple channels in [23, 24, 25] use different modulo thresholds to process the same signal, thereby generating two multi-bit modulo ADC output sample streams. In contrast, our second stream only indicates the position of the folding instances in the first stream. The information carried by this second stream can be represented by a single bit per sample.
@@ II-B Is the 1-bit Folding Information Practical?
One might wonder if generating 1-bit folding information $c[n]$ is practical when implementing a modulo ADC. Early works on modulo ADC hardware have considered recovery methods that utilize the reset count in addition to the self-reset ADC output samples [46, 47]. However, implementing a self-reset ADC with a multi-bit reset count signal may incur a significant increase in power consumption and circuit complexity if the threshold is small [7]. Furthermore, having a multi-bit reset count signal that has the same resolution as the self-reset ADC output samples doubles the storage requirements of the system. Limiting $c[n]$ to 1-bit solves these problems. As shown in Figures 1 and 2, generating $c[n]$ would only require one additional comparator and one adder from the “no $c[n]$” modulo ADC setup. The additional circuitry has minimal impact on the overall power consumption, and its penalty is fixed regardless of $b$. For instance, consider a flash ADC architecture for amplitude quantization. An 8-bit modulo ADC with no $c[n]$ would have $2582^{8}+2=$ 258 comparators [^1] The 2 additional comparators come from the comparators of the folding mechanism. whereas an 8-bit modulo ADC with 1-bit folding information $c[n]$ would have 259 comparators. The difference in comparator counts is fixed regardless of the number of bits for amplitude quantization. We also note that the folding information generated in Figure 1 is one bit less than the folding information generated by the modulo ADC hardware proposed in [32]. In [32], the 2-bit signal indicates three states: ‘no crossing’, ‘crossed $+\lambda^{\prime}$ threshold’, ‘crossed $-\lambda^{\prime}$ threshold’.

### v3

@@ Abstract
The modulo analog-to-digital converter (ADC) is a promising solution to resolve the limited dynamic range (DR) issue of conventional ADCs and achieve an enhanced digital resolution given a fixed quantization bit budget. However, a modulo ADC requires an unfolding scheme to correct the nonlinear distortion introduced by the modulo operation. This paper presents a sliding discrete Fourier Transform (DFT)-based method for fast signal reconstruction given the modulo ADC output sequence and a 1-bit folding information sequence. In contrast to existing DFT-based signal recovery techniques for modulo ADCs, our proposed sliding DFT method reduces the required observation time and minimizes the spectral leakage effects via proper choice of window function parameters. A mean squared error (MSE) performance guarantee is established for the proposed signal recovery algorithm. More precisely, we derive sufficient conditions for the oversampling factor ($\mathrm{OF}$) and the number of quantization bits ($b$) to obtain a specific MSE performance. Our numerical results demonstrate that modulo ADCs equipped with our proposed recovery method can outperform conventional ADCs without modulo for $\mathrm{OF}\geq 4$ and $b\geq 4$. Moreover, the proposed sliding DFT-based recovery method outperforms higher-order difference-based recovery in terms of MSE, particularly in the low-sampling-rate and low-resolution regime. The impact of spectral leakage on the MSE performance of the proposed sliding DFT recovery method is also quantified.
@@ I Introduction
A promising solution to address the DR bottleneck of ADCs is the modulo sampling framework [6, 7]. The idea in modulo sampling is to apply a nonlinear folding operation called modulo to the input signal before the ADC. This modulo pre-processing stage ensures that the signal range is confined within the DR of the ADC. Moreover, for a given amplitude quantization bit budget, modulo ADCs can achieve better digital resolution compared to conventional ADCs as demonstrated in prior works [8, 9].
@@
Since modulo operation intentionally introduces nonlinear distortion to the signal, several studies on modulo sampling have focused on the development of robust recovery techniques to correct the distortion induced by the modulo operation. Reconstruction techniques for modulo sampling include higher-order differences-based approaches [7], discrete Fourier transform (DFT)-based techniques [10, 11, 12, 13, 14, 15], prediction-based filtering [16, 17], iterative methods [18], sparse signal modulo recovery [19, 20, 21, 22], multi-channel modulo structures [23, 24, 25, 26], and thresholding-based schemes[27, 28, 29, 30, 31]. There are also research efforts geared toward the practical hardware implementation of modulo ADCs [10, 32, 33, 34, 35, 36] and their potential applications in radars, communication systems, imaging, and electroencephalogram (EEG) recovery [37, 38, 39, 40, 12, 41, 42, 43, 25, 44, 45, 46].
@@
Another important research direction in modulo ADC is the establishment of theoretical performance guarantees to reconstruct the unfolded signal from modulo samples. In [7], it is shown that the higher-order differences-based reconstruction approach can perfectly reconstruct a finite-energy bandlimited (BL) signal from its modulo samples if the sampling rate used is at least $2\pi e$ ($\approx 17.07$) times the Nyquist rate. This sufficient condition for the sampling rate increases significantly in the presence of bounded noise [7, Theorem 3]. For finite-energy BL signals, perfect reconstruction is possible for any sampling rate approaching the Nyquist rate from above by using a prediction filter with sufficiently long filter length [16]. This approach does not need any side information such as knowledge of unfolded samples and auto-correlation function. However, the proposed prediction filter-based approach did not take into account the impact of finite quantization. Theoretical guarantees for Fourier-based reconstruction had been established in [10]. However, this approach suffers from spectral leakage and may require long observation windows. To avoid spectral leakage, recovery methods based on time-domain thresholding can also be used. Such methods work for generalized modulo sampling models which consider the modulo hysteresis and the folding transients. However, recovery guarantees of thresholding-based recovery fail under ideal modulo nonlinearities. This is explained in Section III-E. The recently developed UNO framework [47] integrates the Unlimited Sampling Framework (USF) [7] with 1-bit quantization, where each folded sample is processed by multiple 1-bit quantizers using varying thresholds. In effect, the set of 1-bit quantizers operates collectively as a multi-bit quantizer. Furthermore, the accompanying theoretical guarantees are expressed through high-probability bounds.
A recent work on modulo ADCs with 1-bit side information [14, 48] established a mean squared error (MSE) performance guarantee when the oversampling factor (OF) is above three and the number of quantization bits is above three. The 1-bit side information identifies the indices in which folding occurred. However, the proofs neglected the spectral leakage effects and the contribution of quantization noise at certain time indices. When the 1-bit folding information is not available in the recovery process, the same MSE guarantees can be achieved by an orthogonal matching pursuit (OMP)-based recovery scheme [48]. However, the sufficient condition for the amplitude quantization bit budget becomes stricter.
@@
In this work, we propose a sliding DFT approach to construct the original input samples from the modulo ADC output samples and an additional 1-bit folding information signal. The recovery procedure uses short-time Fourier transform (STFT) to facilitate successive frame-based unfolding of the modulo samples. To significantly reduce the impact of spectral leakage in STFT computation, a window function with smooth transitions at the edges is multiplied to the frame. A MSE performance guarantee for the proposed recovery method is also established and is compared with the simulated MSE results. Effectively, the use of sliding DFT-based recovery solves the high computational complexity issue of the Fourier-based recovery with 1-bit folding information presented in [14, 48].
@@
  • We compare the MSE of the modulo ADC equipped with our proposed recovery procedure to that of the conventional ADC, i.e., no modulo operation, and that of modulo ADC equipped with higher-order differences (HoD)-based recovery method. Under the established sufficient conditions, we show that the MSE of the modulo ADC which uses the sliding DFT recovery algorithm is in $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ when spectral leakage is neglected and is in $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ otherwise. The parameter $\mathrm{OF}$ is the oversampling factor to be defined in Section II. Meanwhile, the MSE of a conventional ADC is in $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$. We also note that the HoD-based recovery method fails to unfold the modulo ADC output at certain values of $\mathrm{OF}$ when $b=4,5$. Numerical results are presented to demonstrate these behaviors.
Notation: The following notations are used throughout the paper. The sets of real numbers, integers, and natural numbers are denoted as $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{N}$, respectively. When referring to discrete-time signals, the notation $z[n]$ is used to signify $z(nT_{\mathrm{s}})$, assuming that the sampling period $T_{\mathrm{s}}$ is evident from the context. The first-order difference of a discrete-time signal $x[n]$ is denoted as $\underline{x}[n]=x[n]-x[n-1]$, where x[1]=0x[-1]=0 unless stated otherwise. Vectors and matrices are written in bold format (e.g., $\mathbf{z}$, $\mathbf{A}$) while sets are written in calligraphic format (e.g., $\mathcal{S}$). The cardinality of a set $\mathcal{S}$ is written as $|\mathcal{S}|$. If $\mathbf{x}$ is a vectorized form of some discrete-time signal $x[n]$, then $\underline{\mathbf{x}}$ is the vectorized form of $\underline{x}[n]$. The $\ell_{p}$-norm of a vector $\mathbf{x}$ is written as $\|\mathbf{x}\|_{p}$. The set of values of the discrete-time signal $x[n]$ for $n\in\mathcal{S}$ is denoted as $x_{\mathcal{S}}[n]$. The Moore-Penrose inverse of a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ is denoted by $\mathbf{A}^{\dagger}$. We use $\omega$ and $\Omega$ to denote the frequencies for the continuous-time Fourier Transform (CTFT) and discrete-time Fourier Transform (DTFT) spectrum, respectively. We also use the standard Big-O notation $\mathcal{O}(\cdot)$ to describe asymptotic growth rates in this paper.
@@ II-A Signal Acquisition Process
We consider the modulo ADC system shown in Figure 1. The input $f(t)$ is a bandlimited signal with (angular) frequency support $\left[-\frac{\omega_{m}}{2},\;+\frac{\omega_{m}}{2}\right]$. This signal is folded by the modulo operator block to produce the modulo signal $f_{\lambda^{\prime}}(t)$. The operation behind the modulo folding mechanism is shown in Figure 2. This feedback structure for the folding operation is adopted from [34]. Comparator 1 triggers a positive value whenever $f_{\lambda^{\prime}}(t)$ crosses $\lambda$ from below. Similarly, comparator 2 triggers a negative value whenever $f_{\lambda^{\prime}}(t)$ crosses $\lambda$ from above. These trigger signals are fed to a discrete voltage generator (DVG). The voltage level of the DVG is increased (resp. decreased) by $2\lambda^{\prime}$ whenever it receives a positive (resp. negative) trigger signal. A detailed circuit-level implementation of the DVG is presented in [34]. Mathematically, the folded signal can be expressed as
@@
where $\lambda^{\prime}\in(0,\|f(t)\|_{\infty})$ is the modulo threshold. The $\ell_{\infty}$-norm of a waveform is its maximum amplitude. Throughout this paper, we assume ideal modulo nonlinearities. That is, there are no modulo hysteresis and folding transients like in the generalized modulo sampling model presented in [28]. Furthermore, we also assume that the amplitude of the first sample of $f(t)$, i.e. $f[0]$, is within $\lambda^{\prime},\lambda^{\prime}]$. This ensures that no unknown constant offset $2\lambda^{\prime}p$, where $p\in\mathbb{Z}$, is present in the reconstructed signal.
To digitize the modulo signal, $f_{\lambda^{\prime}}(t)$ is first sampled at every $T_{s}$ seconds. The angular sampling rate is $\omega_{s}=\frac{2\pi}{T_{s}}=\mathrm{OF}\times\omega_{m}$, where $\mathrm{OF}\geq 1$ is the oversampling factor. We also define $\rho=\frac{1}{\mathrm{OF}}$. Prior to quantization, a (non-subtractive) dither sequence $d[n]$ is added to $f_{\lambda^{\prime}}[n]$. Samples of the dither sequence are drawn i.i.d. from a triangle distribution with amplitude support $\big{(}-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\big{]}$. The rationale for using triangle dither in the modulo ADC problem setup is explained in Section III-B. The resulting signal is then forwarded to the $b$-bit uniform scalar quantizer $\mathcal{Q}_{b}(\cdot)$ with range $\lambda,+\lambda]$. Consequently, each quantization bin has width $\frac{2\lambda}{2^{b}}$. To prevent the quantizer from being overloaded, the quantizer DR is set to $\lambda=\frac{2^{b}\lambda^{\prime}}{2^{b}-2}$. The quantizer output can be written as
$f_{\lambda^{\prime},\mathrm{q}}[n]=\mathcal{Q}_{b}(f_{\lambda^{\prime}}[n]+d[n])=f[n]+z[n]+\epsilon[n],$ (2)
@@
where $\tau_{k}$ is the location of the level crossings, $g_{1}(t)$ is the pulse shape of the trigger signals, $g_{2}(t)$ is the sampling kernel, and $g(t)=(g_{2}\ast g_{1})(t)$. To simplify the analysis, $g(t)$ is assumed to be a rectangular pulse with width $T_{s}$ and unity amplitude. The signal $c(t)$ is sampled at rate $\omega_{s}$ and then fed to a 1-bit comparator to produce $c[n]$. In essence, $c[n]$ provides information if the signal crossed either $\lambda^{\prime}$ or $\lambda^{\prime}$ in the interval $(nT_{\mathrm{s}},(n+1)T_{\mathrm{s}}]$. We note that even though our modulo ADC generates two channel streams, our setup is different from the multichannel modulo sampling structures presented in [23, 24, 25]. The multiple channels in [23, 24, 25] use different modulo thresholds to process the same signal, thereby generating two multi-bit modulo ADC output sample streams. In contrast, our second stream only indicates the position of the folding instances in the first stream. The information carried by this second stream can be represented by a single bit per sample.
@@ II-B Is the 1-bit Folding Information Practical?
One might wonder if generating 1-bit folding information $c[n]$ is practical when implementing a modulo ADC. Early works on modulo ADC hardware have considered recovery methods that utilize the reset count in addition to the self-reset ADC output samples [49, 50]. However, implementing a self-reset ADC with a multi-bit reset count signal may incur a significant increase in power consumption and circuit complexity if the threshold is small [7]. Furthermore, having a multi-bit reset count signal that has the same resolution as the self-reset ADC output samples doubles the storage requirements of the system. Limiting $c[n]$ to 1-bit solves these problems. As shown in Figures 1 and 2, generating $c[n]$ would only require one additional comparator and one adder from the “no c[n]c[n]” modulo ADC setup. The additional circuitry has minimal impact on the overall power consumption, and its penalty is fixed regardless of $b$. For instance, consider a flash ADC architecture for amplitude quantization. An 8-bit modulo ADC with no $c[n]$ would have $2^{8}+2=258$ comparators [^1] The 2 additional comparators come from the comparators of the folding mechanism. whereas an 8-bit modulo ADC with 1-bit folding information $c[n]$ would have 259 comparators. The difference in comparator counts is fixed regardless of the number of bits for amplitude quantization. We also note that the folding information generated in Figure 1 is one bit less than the folding information generated by the modulo ADC hardware proposed in [32]. In [32], the 2-bit signal indicates three states: ‘no crossing’, ‘crossed $\lambda^{\prime}$ threshold’, ‘crossed $\lambda^{\prime}$ threshold’. Finally, we point out that the modulo ADC with 1-bit folding information has already been implemented in hardware [36], but it follows a different architecture from what is shown in Figures 1 and 2.